OpenStudy (chris215):

Express the sum using sigma notation: 1/2+2/5+3/10+4/17+5/26

OpenStudy (freckles):

top is going up be 1 each time the bottom is a quadratic because the second differences are constant that is 2,5,10,17,26 first differences 3,5,7,9 second difference: 2,2,2

OpenStudy (freckles):

find a,b,c given \[f(n)=an^2+bn+c \text{ and } \\ f(1)=2 \\ f(2)=5 \\ f(3)=10 \\ ,...\]

OpenStudy (freckles):

using just that you will have a system of equations with 3 equations and 3 variables

OpenStudy (freckles):

you can solve that to find a,b, and c

OpenStudy (freckles):

however if you want to try to short cut things you can try to guess the quadratic pattern in the denominator

OpenStudy (chris215):

i got the summation from k = 1 to 5 k+1/(k^2)

OpenStudy (freckles):

here is a hint for the denominator: 2=1^2+1, 5=2^2+1,10=3^2+1,..... also it looks your numerator needs a little fixing